OpenStudy (highschoolmom2010):

Two electrically-charges spheres are suspended from insulated threads a certain distance from each other. There is a certain amount of electrostatic force between them. Describe specifically (not just increase or decrease) what happens to this force in each of the scenarios below: a. The charge on one sphere is reduced by half b. The charge on both spheres is doubled c. The distance between the spheres is increased by a factor of three d. The distance between the sphere is decreased to one-fourth e. The charge of each sphere is doubled and the distance between them is doubled

OpenStudy (anonymous):

I am in an electrical engineering class and i do know that insulated is making it safer to be able to handle the wire so in a way it decreases the voltage of the wires

OpenStudy (michele_laino):

please keep in mind that the electrical force between those two spheres, is given by the subsequent formula: \[\Large F = \frac{1}{{4\pi {\varepsilon _0}}}\frac{{{Q_1}{Q_2}}}{{{d^2}}}\] where Q1, and Q2 are the corresponding electrical charge, and d is their separation

OpenStudy (michele_laino):

charges*